THE PURE ROTATIONAL SPECTRUM OF PIVALOYL CHLORIDE, (CH_{3})_{3}CCOCl, BETWEEN 800 MHz AND 18800 MHz

\underline{GARRY S. GRUBBS II}, CHRISTOPHER T. DEWBERRY, KERRY C. ETCHISION, STEPHEN A. COOKE, Department of Chemistry, The University of North Texas, P. O. Box 305070, Denton, Texas, 76203, U.S.A; MICHEL M. SERAFIN, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Avenue, Charleston, Illinois 61920, U.S.A..

A search accelerated correct intensity Fourier transform microwave (SACI-FTMW) spectrometer has been used to rapidly record the rotational spectrum of pivaloyl chloride between 8 and 18 GHz. A low frequency, cavity-based instrument has then been used to obtain high resolution measurements below 4 GHz. Rotational constants, centrifugal distortion constants and chlorine nuclear quadrupole coupling constants have been determined for both the ^{35}Cl and ^{37}Cl isotopologues. The spectrum provides no evidence concerning internal rotation in the molecule.